OpenStudy (anonymous):

Find the radius and interval of convergence of E n=1 to infinity, of sqr(n)(x^n)

OpenStudy (anonymous):

try ratio test see if it works

OpenStudy (anonymous):

I dont know the steps..

OpenStudy (anonymous):

lim abs[ a(n+1) / a(n) ] n-> inf and we want the limit to be less than 1

OpenStudy (anonymous):

Hmmm okay but what about the problem? Why did you change it?

OpenStudy (anonymous):

lim abs [ sqrt(n+1) x^(n+1) / (sqrt(n) x^n) ] n->inf

OpenStudy (anonymous):

\[\sum_{n=1}^\infty\sqrt n~x^n~~?\]

OpenStudy (anonymous):

I agree with using the ratio test. Figure out the limit: \[\lim_{n\to\infty}\left|\frac{\sqrt{n+1}~x^{n+1}}{\sqrt n~x^n}\right|=L\] The series converges if the limit \(L<1\).